# Efficient Observation Time Window Segmentation for Administrative Data Machine Learning

### 29 Jan 2024 v1

@@
Abstract.
Utilizing administrative data to predict outcomes is an important application area of machine learning, particularly in healthcare. Most administrative data records are timestamped and the pattern of records over time is a key input for machine learning models. This paper explores how best to divide the observation window of a machine learning model into time segments or “bins”. A computationally efficient process is presented that identifies which data features benefit most from smaller, higher resolution time segments. Results generated on healthcare and housing/homelessness administrative data demonstrate that optimizing the time bin size of these high priority features while using a single time bin for the other features achieves machine learning models that are simpler and quicker to train. This approach also achieves similar and sometimes better performance than more complex models that default to representing all data features with the same time resolution.
@@ Abstract.
administrative data, time window segmentation, machine learning, hospital, homelessness
1. Introduction
@@
The collection of administrative data is now commonplace. While definitions vary, administrative data is most generally though of as “information collected by government or private sector organizations as part of their ongoing operations” (statisticscanada2023, 2). This data is typically collected as part of the delivery of a service that can range from interaction with the justice system, filing taxes, progressing through the education system or accessing housing and homelessness services (lyon2015, 23, 42, 6, 17). However, with the extensive use of electronic medical records (EMRs), it is arguable that healthcare is the one of the most important sectors to make use of administrative data (mazzali2015, 25).
One of the most exciting emerging applications of administrative data is combining it with machine learning to predict outcomes both at the individual and population levels. This area of research is particularly active in the healthcare domain (morid2023, 28, 11) and is increasingly being applied to the social determinants of health such as housing and homelessness (vanberlo2021, 41, 40).
@@
The first contribution of this paper is to present a detailed examination of how administrative data should be pre-processed in the time domain in order to produce the best possible ML performance. Since most administrative data entries are time stamped, administrative data are fundamentally time series data. The time dimension carries important trend information like whether a patient’s condition in hospital is improving or getting worse with time. The length of time between events is also important since there is clearly a difference between someone who accesses an emergency housing shelter seven times in one week versus seven times in one year. The importance of the time dimension is not lost on most machine learning researchers. For example, the excellent survey by Morid, et. al. (morid2023, 28) summarizes over 76 papers applying machine learning to medical administrative data and distinguishes them in part by how they account for the temporal nature of that data.
While the temporal nature of administrative data is recognized as important, there is little consensus regarding how the time dimension should be represented. A popular approach in (morid2023, 28) is to use time information to create an ordered sequence of the events that fall within an observation window. However, a large proportion instead take the temporal matrix approach where temporal data is represented as a matrix. In this matrix, rows correspond to specific events or data features and columns correspond to regularly spaced time intervals within the observation window.
This paper will focus on the temporal matrix framework. While sequence representation is efficient, the temporal matrix approach has the potential to achieve better model performance. This is based in part on the notion that representing when an event does not happen can often be as valuable as representing when it does (lipton2016, 21). In this paper, the regularly spaced time intervals in the temporal matrix will be referred to as the observation window time bins. When dealing with countable events, each cell in the temporal matrix contains the number of times a particular event occurs within the corresponding time bin. The observation window used by traditional survival analysis would correspond to a single time bin represented by a matrix with a single column (klein2006, 16). At the other extreme, signal data from a biomedical sensor would be a single row matrix with columns representing time bins with width equal to the inverse of the Nyquist sampling rate (proakis1995, 29).
The second contribution of this paper is to present a framework that can be used to determine the temporal matrix time bin size that maximizes machine learning model performance for a given observation window. We demonstrate that time bin size behaves similar to any other machine learning model hyperparameter where it can contribute to the over or under fitting of the model to the data. We will show that simply increasing the number of time bins (ie. reducing bin size) for all data features is not a good strategy since more bins can greatly increase the complexity of the machine learning model. This leads to increased training time and a higher risk of overfitting the data. Consistent with (svenson2020, 38), we demonstrate that not all data features require bins with the same time resolution. We build on this idea and present a computationally efficient method to pre-process data features to determine which ones will benefit from increased time resolution and which ones will not. Our results will demonstrate that this method leads to machine learning models that are efficient to train and perform as well or better than models that represent all features with the same time resolution.
2. Related Work
@@
There is a wide variation in how the temporal matrix framework is applied in practice. A number of studies adopt the single time bin approach where the start of the window coincides with a patient first appearing in the data (el-rashidy2020, 31, 30, 11). A single time bin fits within the temporal matrix framework but clearly does not allow the machine learning model to benefit of trends in data features over time. Several other studies demonstrate the feasibility of working with regularly spaced time bins but do not justify their choice of bin size (lee2015, 20, 21, 11, 43, 39, 27).
@@
There has been some work on demonstrating how the temporal matrix time bin size affects machine learning model performance. Varying the size of a single bin that starts from time zero has been investigated in (el-rashidy2020, 31, 30). Min, et. al. varies the size of multiple uniformly sized time bins and demonstrates improved performance with larger bins (min2019, 27). Svenson, et. al. also investigate different time bin resolutions and demonstrate that not all data features require the same bin size in order to maximize model performance (svenson2020, 38).
To date, there has not been a rigorous investigation of how best to determine time bin size or an exploration of how time bin size affects machine learning model complexity in general. Using a methodical approach to adjusting time bin size, our paper will demonstrate that increasing the number of time bins contributes to machine learning model complexity in a way similar to other hyper-parameters like decision tree height or the number of neural network layers. It is well understood that an overly complex machine learning model risks data overfitting and can be resource intensive to train (hastie2009, 12). We will demonstrate that an excessive number of time bins causes both of these two pitfalls to occur.
To the author’s knowledge, only (svenson2020, 38) has investigated the notion that not all data features require the same time bin resolution. While cycling through the range of all possible bin size combinations for all data features as part of a hyperparameter tuning grid search is theoretically possible, it is often not computationally feasible. Our paper will propose a computationally efficient technique called time-series analysis to investigate binning (TAIB) that identifies which data features will benefit from increased time resolution. TAIB makes use of the dynamic time warping (DTW) technique (sakoe1978, 33) that has been used in many applications for robustly comparing the difference between time sequences.
3. Data Sets
@@
The first administrative data set employed in this paper is MIMIC-III (Medical Information Mart for Intensive Care III), a comprehensive and publicly available database widely used in critical care research (johnson2016, 13). Our objective with the MIMIC-III data set is to predict patient mortality in the intensive care unit (ICU), a significant challenge in medical analytics (liu2022predicting, 22, 32). This task is formulated as a binary classification problem, where we use data from the first 48 hours of a patient’s ICU stay to predict mortality.
@@
The MIMIC-III database encompasses a wide array of healthcare information across 26 tables, including detailed records of patient demographics, vital signs, laboratory test results, and more. For our mortality prediction analysis, we particularly focus on the CHARTEVENTS and LABEVENTS tables. These tables are crucial as they contain physiological metrics and lab results that are integral to assessing a patient’s condition. As in (purushotham2018, 30), the specific features selected are those related to the SAPS II (Simplified Acute Physiology Score II) score (legall1993, 18). SAPS II is a globally recognized severity-of-disease classification system, designed to evaluate the risk of ICU patient mortality. The selected features include age, vital sign information (heart rate, systolic and diastolic blood pressure, and body temperature) and laboratory results (blood urea nitrogen levels, white blood cell count, potassium levels, sodium levels, bicarbonate levels, bilirubin levels, and Glasgow Coma Scale). Consistent with (purushotham2018, 30), the MIMIC-III data is filtered to include only first ICU admissions. We exclude individuals who left the ICU within the first 48 hours and those under 15 years old.
@@
Table 1. DI data set record entry types.
The DI data set also contains free text case notes associated with most entries. To protect client privacy, these fields were replaced with keyword count data features by DI staff prior to its release. The keyword categories, the number of search keywords in each category and some examples of those search words are provided in in Table 2. For example, if a record contained the free text comment string ”The client was involved in a fight and was punched. Police were called.”, that comment field would be replaced by keyword counts of 2 for the Conflict category, 1 for Police/Justice and 0 for all other categories.
@@
Table 2. DI data set comment field keyword categories.
@@
Following established practice for homelessness administrative data analysis (kuhn1998, 17), a clustering algorithm is applied to the historical shelter use patterns of people present in the DI data to label each person as having a chronic, episodic or transitional pattern of shelter use. Chronic shelter use corresponds to long term, sustained shelter use often over many years. Episodic patterns are also longer term but much more sporadic. Finally, the transitional label is assigned to people who use shelter for only a short period of time. Chronic shelter users are typically a minority of the people accessing emergency shelter but their sustained use means they consume the majority of system resources (kuhn1998, 17). As a result, they are often prioritized for supportive housing programs.
@@
Our objective with the DI data is to use the first 90 days of a person’s record of shelter use to predict whether that person will become a chronic emergency shelter user (the positive case) or an episodic or transitional user (the negative case). Cluster analysis is typically only performed on historical data to determine a person’s shelter use patterns in retrospect. A machine learning method that uses a person’s first 90 days in shelter to predict if they are at risk of becoming a chronic shelter user could be used in real time to connect that person to supports.
4. Methods
4.1. Data Representation
@@
Assume a data set containing records for $N$ people. Each person is associated with one or more time stamped records and each record contains $K$ data features in addition to the time stamp and person identification field. Following (morid2023, 28), this data can be represented by $N$ temporal matrices, each matrix representing the experience of one person. While each person will typically have a different number of records, the $N$ temporal matrices can be made the same size by defining an observation window of length $T_{O}$ relative to each person’s earliest record. The observation window can then be divided into $L=T_{O}/T_{B}$ equally spaced time bins, where $T_{B}$ is the duration of each bin. Bins where no event occurs for a particular person can be filled with zeros or null values. The data set therefore contains $N$ instances where each instance is an $L\times K$ temporal matrix associated with one person.
@@
Categorical data features are assumed to be one hot encoded so that each categorical value is a separate data feature. A temporal matrix cell for one of these categorical features contains the number of times that categorical value occurs during the cell’s corresponding time bin. Continuous data features are represented in this paper by taking the average of all the continuous values that fall into each bin. However, in general, this averaging operation could be replaced by any mathematical operation on the continuous values. Continuous value data feature bins that contain no time stamped values are assigned a null value. Static data features are copied across all time bins.
Recurrent neural networks (RNNs) and their variants are designed for multi-dimensional data features that occur at regularly spaced time intervals. This means that these algorithms are restricted to operating on data in the temporal matrix representation described above. However, machine learning algorithms not designed specifically for regularly sampled time domain data are not constrained by the requirement that each of the $K$ data features be represented using the same number of time bins.
@@
Figure 1. Temporal and feature matrix, single user, $K=3$, $T_{O}$ = 6 hr.
The feature matrix format can be used by any machine learning algorithm not designed specifically for regularly spaced temporal values. These algorithms include all of the “classic” machine learning models (ie. logistic regression, variations on the decision tree, support vector machines) (hastie2009, 12) as well as multi-layer perceptron and convolutional neural networks (Goodfellow-et-al-2016, 8). These algorithms do not depend on a common temporal bin size since they treat each of the $V$ values in a feature matrix row as distinct data features. However, these algorithms can still benefit from temporal trends that help to differentiate people. For example, a person accessing an emergency room sporadically will have a different pattern of values in their feature vector than someone else who accesses an emergency room regularly. Any machine learning algorithm benefits from different data feature patterns, even if the algorithm isn’t explicitly designed to recognize that the patterns vary over time.
4.2. Efficiently Assigning Bin Size
The size of the time bins used in the temporal matrix and feature matrix representations described in Section 4.1 is an important design parameter. At one extreme, a single bin (ie. $T_{B,k}=T_{O}$, $k=1\ldots K$) results in both the temporal and feature matrix representations reducing to $N$ vectors of length $K$. This reduces data set size and decreases model training time. However, it also discards all temporal information which degrades model performance for many data sets. At the other extreme, $T_{B,k}$ could be set to the maximum resolution available, often 1 second for many health data sets. While this captures all temporal trends in the data, this may also result with a large and very sparse data set with many null entries. The penalty is greatly increased training time and degraded performance due to the curse of dimensionality (hastie2009, 12).
@@ 4.2. Efficiently Assigning Bin Size
(1) $S_{k,L_{k}}=\frac{1}{|\mathcal{N}|}\sum_{n\in\mathcal{N}}{\rm DTW}\left(n,C_{\mathcal{P}}\right)+\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}{\rm DTW}\left(p,C_{\mathcal{N}}\right)$
where ${\rm DTW}(a,b)$ is the dynamic time warped Euclidean distance between vectors $a$ and $b$ (sakoe1978, 33). This process is repeated for each value of $L_{k}$ in the simply ordered set $\mathcal{L}$. The corresponding values of $S_{k,L_{k}}$ calculated for each value in $\mathcal{L}$ form a $|\mathcal{L}|$ element vector $\mathbf{S}_{k}$.
Dynamic time warping (DTW) (sakoe1978dynamic, 34) warps the time axes of two vectors $a$ and $b$ in order to achieve the maximum co-incidence between them (han2011data, 10). Since the objective is to spot when adjusting $L_{k}$ results in a noticeable difference between features, DTW can be seen as a conservative choice since the warping is an attempt to minimize the distance between two feature vectors. This means that when a difference between vectors is observed, using DTW distance allows us to be more confident it is a meaningful change caused by adjusting $L_{k}$ and not a random artifact of the data.

### 12 Mar 2024 v2

@@
Abstract
Machine learning models benefit when allowed to learn from temporal trends in time-stamped administrative data. These trends can be represented by dividing a model’s observation window into time segments or bins. Model training time and performance can be improved by representing each feature with a different time resolution. However, this causes the time bin size hyperparameter search space to grow exponentially with the number of features. The contribution of this paper is to propose a computationally efficient time series analysis to investigate binning (TAIB) technique that determines which subset of data features benefit the most from time bin size hyperparameter tuning. This technique is demonstrated using hospital and housing/homelessness administrative data sets. The results show that TAIB leads to models that are not only more efficient to train but can perform better than models that default to representing all features with the same time bin size.
@@ Abstract
Keywords administrative data, time window segmentation, machine learning, hospital, homelessness
1 Introduction
@@
Utilizing data features from administrative data to predict outcomes for individuals or groups is an exciting application of machine learning. Administrative data typically consist of records collected while delivering services that can include the justice system, filing taxes, progressing through the education system or accessing housing and homelessness services [22, 38, 6, 16]. However, with the extensive use of electronic medical records (EMRs), it is arguable that healthcare is the one of the most important sectors to make use of administrative data [24].
@@
Since most administrative data entries are time stamped, administrative data are fundamentally time series data. The time dimension carries important trend information for a machine learning model like whether a patient’s condition in hospital is improving or getting worse with time. The length of time between events is also important since there is clearly a difference between someone who accesses an emergency housing shelter seven times in one week versus seven times in one year. The importance of the time dimension is not lost on most machine learning researchers. For example, the excellent survey by Morid, et. al. [27] summarizes over 76 papers applying machine learning to medical administrative data and distinguishes them in part by how they account for the temporal nature of that data.
While a popular approach is to simply use temporal information to organize data features into an ordered sequence, understanding when an event does not occur can be as valuable as representing as when it does [20]. As a result, a large number of studies represent temporal administrative data using the temporal matrix approach [27]. In a temporal matrix, columns correspond to specific events or data features and rows correspond to regularly spaced time intervals within the observation window. These time intervals are referred to as time bins.
Time bin duration or “size” is an important temporal matrix parameter. While the simplistic approach is to use the same bin size for all data features, it has been demonstrated that performance and model complexity can be improved when representing different features using different time resolutions [36, 1]. However, the challenge is that allowing each feature to have its own bin size means that the model’s hyperparameter search space will grow exponentially with the number of data features.
The contribution of this paper is to present a pre-processing technique that determines which data features benefit the most from time bin size tuning and which features can be represented using a single time bin. Using dynamic time warping [31], features are ranked according to a new metric that quantifies how much their power for discrimininating between data set labels improves as they are represented at a higher time resolution. A handful of features from the top of this list are selected for time bin tuning and the remaining features are represented using a single time bin. Using data sets from the medical and housing/homelessness domains, we will demonstrate that this approach results in simpler models that are quicker to train while still performing as well or better than when all data features are represented using the same time resolution.
2 Related Work
@@
There is a wide variation in how the temporal matrix framework is applied in practice. A number of studies use a single time bin where the start of the window coincides with a patient first appearing in the data [29, 28, 11]. A single time bin fits within the temporal matrix framework but clearly does not allow the machine learning model to benefit from the trends that appear in data features over time. Several other studies use multiple time bins and apply the same bin size to all data features but do not justify their choice of bin size [19, 20, 11, 39, 37, 26].
@@
There has been some work on demonstrating how the temporal matrix time bin size affects machine learning model performance. Varying the size of a single bin that starts from time zero has been investigated in [29, 28]. Min, et. al. varies the size of multiple uniformly sized time bins and demonstrates improved performance with larger bins [26].
A limited number of studies have also investigated using different bin sizes for different features. Svenson, et. al. investigate different time bin resolutions and demonstrate that not all data features require the same bin size in order to maximize model performance [36]. Barzegar, et. al. take a multi-rate signal processing approach where they determine how best to represent different data features using different sample rates [1].
Most machine learning models can accomodate data features represented using different time resolutions. Classical machine learning models and non-recurrent neural networks simply treat each time bin of each data feature as a separate data feature input. Nothing in the machine learning model structure requires that these bins are the same size. The traditional forms of recurrent neural network (RNN’s) models do assume that all data features are represented as a multi-dimensional input with a common time resolution. However, RNN variations that include ensemble RNNs [1] and channelized long short-term memory (LSTM) networks [11] have been proposed that allow for different features to be represented with different time resolutions.
Time bin size must be selected with care and not all data features need to be represented using the same time resolution [36, 1]. Choosing a common time bin size that is uncessarily small for some data features results in an overly complex model. This complexity risks data overfitting and can be resource intensive to train [12]. However, the size of an exhaustive search for a unique bin size for each data feature grows exponentially with the number of data features and is impractical for many applications.
To the authors’ knowledge, only [1] has proposed a non-exhaustive search approach for selecting different time resolutions for each feature. However, the method proposed by [1] is based on physics informed machine learning for vehicle sensing applications and is not suitable for general administrative data applications.
This paper addresses this gap by proposing a pre-processing technique called time-series analysis to investigate binning (TAIB) that is suitable for generic administrative data applications. TAIB produces a ranking of data features in decreasing order of how much the machine learning model will benefit from representing that feature using multiple time bins with increasing resolution. TAIB makes use of the dynamic time warping (DTW) technique [31] that has been used in many applications for robustly comparing the difference between time sequences.
3 Data Sets
@@
The first administrative data set employed in this paper is MIMIC-III (Medical Information Mart for Intensive Care III), a comprehensive and publicly available database widely used in critical care research [13]. Our objective with the MIMIC-III data set is to predict patient mortality in the intensive care unit (ICU), a significant challenge in medical analytics [21, 30]. This task is formulated as a binary classification problem, where we use data from the first 48 hours of a patient’s ICU stay to predict mortality.
@@
The MIMIC-III database encompasses a wide array of healthcare information across 26 tables, including detailed records of patient demographics, vital signs, laboratory test results, and more. For our mortality prediction analysis, we particularly focus on the CHARTEVENTS and LABEVENTS tables. These tables are crucial as they contain physiological metrics and lab results that are integral to assessing a patient’s condition. As in [28], the specific features selected are those related to the SAPS II (Simplified Acute Physiology Score II) score [17]. SAPS II is a globally recognized severity-of-disease classification system, designed to evaluate the risk of ICU patient mortality. The selected features include age, vital sign information (heart rate, systolic and diastolic blood pressure, and body temperature) and laboratory results (blood urea nitrogen levels, white blood cell count, potassium levels, sodium levels, bicarbonate levels, bilirubin levels, and Glasgow Coma Scale). Consistent with [28], the MIMIC-III data is filtered to include only first ICU admissions. We exclude individuals who left the ICU within the first 48 hours and those under 15 years old.
@@
Table 1: DI data set record entry types.
The DI data set also contains free text case notes associated with most entries. To protect client privacy, these fields were replaced with keyword count data features by DI staff prior to the release of the data. The keyword categories, the number of search keywords in each category and some examples of those search words are provided in in Table 2. For example, if a record contained the free text comment string “The client was involved in a fight and was punched. Police were called.”, that comment field would be replaced by keyword counts of 2 for the Conflict category, 1 for Police/Justice and 0 for all other categories.
@@
Table 2: DI data set comment field keyword categories.
@@
Following established practice for homelessness administrative data analysis [16], a clustering algorithm is applied to the historical shelter use patterns of people present in the DI data to label each person as having a chronic, episodic or transitional pattern of shelter use. Chronic shelter use corresponds to long term, sustained shelter use often over many years. Episodic patterns are also longer term but much more sporadic. Finally, the transitional label is assigned to people who use shelter for only a short period of time. Chronic shelter users are typically a minority of the people accessing emergency shelter but their sustained use means they consume the majority of system resources [16]. As a result, they are often prioritized for supportive housing programs.
@@
Our objective with the DI data is to use the first 90 days of a person’s record of shelter use to predict whether that person will become a chronic emergency shelter user (the positive case) or an episodic or transitional user (the negative case). Cluster analysis is typically only performed on historical data to determine a person’s shelter use patterns in retrospect. A machine learning method that uses a person’s first 90 days in shelter to predict if they are at risk of becoming a chronic shelter user could be used in real time to connect that person to support programs.
4 Methods
4.1 Data Representation
@@
Assume a data set containing records for $N$ people. Each person is associated with one or more time stamped records and each record contains $K$ data features in addition to the time stamp and person identification field. Following [27], this data can be represented by $N$ temporal matrices, each matrix representing the experience of one person. While each person will typically have a different number of records, the $N$ temporal matrices can be made the same size by defining an observation window of length $T_{O}$ relative to each person’s earliest record. The observation window can then be divided into $L=T_{O}/T_{B}$ equally spaced time bins, where $T_{B}$ is the duration of each bin. Bins where no event occurs for a particular person can be filled with zeros or null values. The data set therefore contains $N$ instances where each instance is an $L\times K$ temporal matrix associated with one person.
@@
Categorical data features are assumed to be one hot encoded so that each categorical value is a separate data feature. A temporal matrix cell for one of these categorical features contains the number of times that categorical value occurs during the cell’s corresponding time bin. Continuous data features are represented in this paper by taking the average of all the continuous values that fall into each bin. In general, this averaging operation could be replaced by any mathematical operation on the continuous values. Continuous value data feature bins that contain no time stamped values are assigned a null value. Static data features are copied across all time bins.
As noted in Section 2, traditional recurrent neural networks are designed for multi-dimensional data features that occur at regularly spaced time intervals. This is consistent with the temporal matrix representation described above. To accomodate different features with different time resolutions, temporal matrix representation can be generalized.
@@
Figure 1: Temporal and feature matrix, single user, $K=3$, $T_{O}$ = 6 hr.
The feature matrix format can be used by any machine learning algorithm not designed specifically for regularly spaced temporal values. These algorithms include all of the “classic” machine learning models (ie. logistic regression, variations on the decision tree, support vector machines) [12] as well as multi-layer perceptron and convolutional neural networks [8]. These algorithms do not depend on a common temporal bin size since they treat each of the $V$ values in a feature matrix row as distinct data features. However, these algorithms can still benefit from temporal trends that help to differentiate people. For example, a person accessing an emergency room sporadically will have a different pattern of values in their feature vector than someone else who accesses an emergency room regularly. Any machine learning algorithm benefits from different data feature patterns, even if the algorithm isn’t explicitly designed to recognize that the patterns vary over time.
4.2 Efficiently Assigning Bin Size
The size of the time bins used in the temporal matrix and feature matrix representations described in Section 4.1 is an important design parameter. At one extreme, a single bin (ie. $T_{B,k}=T_{O}$, $k=1\ldots K$) results in both the temporal and feature matrix representations reducing to $N$ vectors of length $K$. This reduces data set size and decreases model training time. However, it also discards all temporal information which degrades model performance for many data sets. At the other extreme, $T_{B,k}$ could be set to the maximum resolution available, often 1 second for many health data sets. While this captures all temporal trends in the data, this may also result with a large and very sparse data set with many null entries. The penalty is greatly increased training time and degraded performance due to the curse of dimensionality [12].
@@ 4.2 Efficiently Assigning Bin Size
$S_{k,L_{k}}=\frac{1}{|\mathcal{N}|}\sum_{n\in\mathcal{N}}{\rm DTW}\left(n,C_{\mathcal{P}}\right)+\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}{\rm DTW}\left(p,C_{\mathcal{N}}\right)$ (1)
where ${\rm DTW}(a,b)$ is the dynamic time warped Euclidean distance between vectors $a$ and $b$ [31]. This process is repeated for each value of $L_{k}$ in the simply ordered set $\mathcal{L}$. The corresponding values of $S_{k,L_{k}}$ calculated for each value in $\mathcal{L}$ form an $|\mathcal{L}|$ element vector $\mathbf{S}_{k}$.
Dynamic time warping (DTW) [32] warps the time axes of two vectors $a$ and $b$ in order to achieve the maximum co-incidence between them [10]. Since the objective is to spot when adjusting $L_{k}$ results in a noticeable difference between features, DTW can be seen as a conservative choice since the warping is an attempt to minimize the distance between two feature vectors. This means that when a difference between vectors is observed, using DTW distance allows us to be more confident it is a meaningful change caused by adjusting $L_{k}$ and not a random artifact of the data.